OpenStudy (anonymous):

2 2+4 2+4 2+4+6 2+4+6+8 2+4+6+8+10 Look at the table. Make a conjecture about the sum of the first 30 positive even numbers. The sum is 29 × 30. The sum is 30 × 31. The sum is 30 × 30. The sum is 31 × 32.

OpenStudy (anonymous):

1: 2 = 2*1 2: 2 + 4 = 6 = 3*2 3: 2 + 4 + 6 = 12 = 4* 3 ... 30: 31*30

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no problem